OpenStudy (anonymous):

Let f(x)=x^2+2x-2 and g(x)=2-3x. Find the following: a.) f(-1) b.) g(-3) c.) f(a+2)

OpenStudy (lalaly):

f(-1) =(-1)^2 +2(-1)-2 = 1-2-2 = -3 g(-3) = 2-3(3) = 2-9 = -7 f(a+2) = (a+2)^2 +2(a+2)-2

OpenStudy (anonymous):

thank you @lalaly :)

OpenStudy (anonymous):

also note that \[f(a+h)=(a+2)^2+2(a+2)-2\] \[=a^2+4a+4+2a+4-2\] \[=a^2+6a+6\]

OpenStudy (lalaly):

yesss:D

OpenStudy (anonymous):

Okay, thank you @satellite :)

OpenStudy (anonymous):

lala, i have a question

OpenStudy (anonymous):

g(-3) = 2-3(3) = 2-9 = -7 isn't it supposed to be 11? because 2-3(-3) = 2+9